# Hybrid thermalization in the large $N$ limit

### v1

@@ 1 Introduction
One such approach is the semi-holographic formalism Faulkner:2010tq ; Mukhopadhyay:2013dqa ; Iancu:2014ava ; Doucot:2020fvy ; Doucot:2024hzq which provides such a unified and consistent framework. In this framework, one considers two subsectors described by perturbative and holographic approaches, respectively. These two sectors are coupled via their individual effective metrics in a way that the full energy-momentum tensor is manifestly local and is conserved in the physical background metric (which is Minkowski space for practical purposes). This democratic coupling scheme, proposed in Banerjee:2017ozx ; Kurkela:2018dku and reviewed in Sec. 2, can be stated in terms of an effective action and is consistent with the Wilsonian renormalization group. In the large $N$ limit, it has been shown that the dynamics of the full system can be solved explicitly in a self-consistent manner Mukhopadhyay:2015smb ; Ecker:2018ucc ; Mondkar:2021qsf .
@@
The phenomenological implications of the semi-holographic approach can be derived primarily from the hybrid hydrodynamic attractor surface which was studied explicitly in Mitra:2020mei ; Mitra:2022uhv in the context of the Bjorken flow by approximating the dynamical description of each subsector in terms of a Müller-Israel-Stewart (MIS) formulation Muller:1967zza ; Israel:1979wp ; Baier:2007ix . It was shown that not only on the attractor surface, but also for phenomenologically relevant flows (evolving towards the attractor surface) which can be matched to glasma-type approaches Kowalski:2007rw at relevant proper times, the bottom-up thermalization scenario Baier:2000sb is realized (i.e. it can be demonstrated analytically that the energy density of the hard sector dominates over that of the soft sector) while the full system can be described in terms of a single fluid at late time. Furthermore, the comparison of hydrodynamization of the individual subsectors leads to novel insights into small system collisions. It was shown that although the perturbative sector hydrodynamizes later in Pb-Pb type collisions compared to p-p type collisions, the soft sector can hydrodynamize earlier in p-p type collisions compared to Pb-Pb type collisions, both on the hybrid attractor surface and in phenomenologically relevant flows Mitra:2022uhv . This follows from how the hydrodynamization times scale with the total energy density in both subsectors.
@@
In this work, we investigate fundamental issues about thermalization in the semi-holographic approach. The two subsystems are isolated in their effective metrics, although they do exchange energy and momentum from the point of view of the physical background metric. In the large $N$ limit, the subsystems also have their individual entropy currents. This implies that, for generic initial conditions, the full system can relax to pseudo-equilibrium states, where each subsystem has a distinct physical temperature, without reaching global equilibrium. In fact, in Sec. 5, we explicitly demonstrate that the latter can indeed be the case. Furthermore, in the democratic coupling scheme, it is not obvious a priori that the following are satisfied:
@@
In Kurkela:2018dku , conditions (i) and (ii) were shown to hold generally in the semi-holographic democratic coupling scheme. We review these proofs in Sec. 3 while stating the issue of pseudo-equilibrium states explicitly and extending the concept of statistical entropy to the latter. In Sec. 4 we prove that the condition (iii) mentioned above is indeed satisfied, i.e., the global equilibrium state, which can be consistently defined, is indeed the unique maximum entropy state of the micro-canonical ensemble. In Sec. 5, we employ the coarse-grained MIS description for both sectors and demonstrate that when the average total energy density is large, a typical non-equilibrium state of the full isolated system does indeed relax to the global equilibrium state, by formulating this criterion in precise mathematical terms. Finally, in Sec. 6, we compare our results with quantum statistical approaches such as the eigenstate thermalization hypothesis, and discuss further studies which are necessary for phenomenological applications.
Figure 1: The isolated system, $\mathcal{F}$, is composed of two subsystems, $\mathcal{U}$ and $\tilde{\mathcal{U}}$, which individually satisfy their own conservation equations eq. 4 in their respective background metric. The two subsystems are coupled and interact via deformations of their metric, given by sec. 2.1.
@@ 2 Semi-holography
The semi-holographic framework proposes a democratic formulation Banerjee:2017ozx in which at any energy scale, the full theory is defined by both the weakly coupled perturbative sectors as well as the strongly coupled non-perturbative sector, which has a classical gravity dual in the large N limit. The two sectors, otherwise isolated from each other, are allowed to interact by deforming the marginal and relevant couplings of each sector, which get promoted to local algebraic functions of the corresponding operators, and this can be implemented via an effective action Banerjee:2017ozx ; Kurkela:2018dku . This coupling scheme ensures the renormalizability of the effective action of the full system in the large N limit. The democratic coupling also allows for the construction of low-energy dynamics of the full system from the effective low-energy descriptions of the individual sector. The two sectors living in their respective effective metric background are assumed to share the same topological space so that we can describe the full system, including these sectors, using the same coordinates. Moreover, the energy-momentum tensor of the full system is conserved in the physical background while the subsectors are closed in their respective background metric and are allowed to exchange energy and momentum with respect to the physical background metric. This picture is summarized in Fig. 1.
@@ 2.1 Effective metric coupling
Consider a dynamical system $\mathcal{F}$ in a fixed metric background $g^{(B)}_{\mu\nu}$. The physical system is composed of two sub-sectors $\mathcal{U}$ and $\tilde{\mathcal{U}}$, which correspond to the perturbative and non-perturbative degrees of freedom, see Fig. 1. The two sub-sectors live in their respective metric background $g_{\mu\nu}$ and $\tilde{g}_{\mu\nu}$. The effective action of the system that captures the deformation of the marginal/relevant coupling of the individual sectors and their interaction is Kurkela:2018dku
@@
in $d$-dimensional spacetime. ${S}$ and $\tilde{{S}}$ are the effective actions of the respective sectors, $\mathcal{U}$ and $\tilde{\mathcal{U}}$ and $\phi$ and $\tilde{\phi}$ denote all the matter fields in the two sub-sectors. $\gamma$ and $\gamma^{\prime}$ are the dimensionful semi-holographic coupling of mass dimension $-d$ whose ratio gives the dimensionless coupling $r:=-\frac{\gamma^{\prime}}{\gamma}$. The effective metrics $g_{\mu\nu}$ and $\tilde{g}_{\mu\nu}$ act as auxiliary fields and capture the marginal deformation of the respective sectors. As mentioned earlier, the deformation is taken care of by promoting the effective metrics as local algebraic functions of the operators of the sub-sectors. The algebraic equations of the effective metric are obtained by varying the action (1) with respect to the individual effective metrics. The lowest-order metric coupling equations that dictate the coupling between the sub-sectors explicitly are
@@
$T^{{\mu\nu}}$ ::=:= $-\frac{2}{\sqrt{-g^{(B)}}}\frac{\delta{S}_{tot}}{\delta g^{(B)}_{\mu\nu}},$
@@
$\Delta K=$displaystyle-- $\frac{\gamma}{2}~{}\Bigg{(}t^{\rho\alpha}\frac{\sqrt{-g}}{\sqrt{-g^{(B)}}}\Bigg{)}~{}g_{\alpha\beta}^{(B)}~{}\Bigg{(}\tilde{t}^{\beta\sigma}\frac{\sqrt{-\tilde{g}}}{\sqrt{-g^{(B)}}}\Bigg{)}~{}g^{(B)}_{\sigma\rho}$ (6)
@@
As the full action is not invariant under individual diffeomorphisms of the subsystems but only under their diagonal combination, we can expect that the Ward identities of the individual subsystems should imply the Ward identity for the full system.
@@ 3.1 The global equilibrium
We can readily see that given the (constant) physical temperature $\mathcal{T}_{0}$ and the equations of states of the two subsystems ($P(\varepsilon)$ and $\tilde{P}(\tilde{\varepsilon})$), we can determine the energy-momentum tensor of the full system (for an illustration, see next subsection), and thus the total energy density $\mathcal{E}$ and the total pressure $P$. The local temperature $\mathcal{T}$ of the full system and its entropy density $\mathcal{S}$ can then be obtained from the thermodynamic identities
@@
where $\varepsilon(\tilde{\varepsilon})$ and $P(\tilde{P})$ are the energy density and pressure of the respective subsystems. The question is whether $\mathcal{T}$ and $\mathcal{S}$ determined via (13) agree with (11) and (12) as required by consistency with basic principles of thermodynamics and statistical mechanics.
@@ 3.2 Proof of thermodynamic/statistical consistencies
[…]
@@ 3.3 The persistence of pseudo-equilibrium
However, there is another problem of dynamical origin. More generally, if the two subsystems have entropy currents $s_{1}^{\mu}$ and $s_{2}^{\mu}$ (which in equilibrium take the forms $s_{1}^{\mu}=s_{1}u^{\mu}$ and $s_{2}^{\mu}=s_{2}\tilde{u}^{\mu}$) that satisfy $\nabla_{\mu}s_{1}^{\mu}\geq 0$ and ~$\tilde{\nabla}_{\mu}s_{2}^{\mu}\geq 0$ in their respective backgrounds, then it is easy to see that the total entropy current $\mathcal{S}^{\mu}$ can be defined as

### v2

@@ 1 Introduction
One such approach is the semi-holographic formalism Faulkner:2010tq ; Mukhopadhyay:2013dqa ; Iancu:2014ava ; Doucot:2020fvy ; Doucot:2024hzq which provides such a unified and consistent framework. In this framework, one considers two subsectors described by perturbative and holographic approaches, respectively. These two sectors are coupled via their individual effective metrics in a way that the full energy-momentum tensor is manifestly local and is conserved in the physical background metric (which is Minkowski space for practical purposes). This democratic coupling scheme, proposed in Banerjee:2017ozx ; Kurkela:2018dku and reviewed in Sec.˜2, can be stated in terms of an effective action and is consistent with the Wilsonian renormalization group. In the large $N$ limit, it has been shown that the dynamics of the full system can be solved explicitly in a self-consistent manner Mukhopadhyay:2015smb ; Ecker:2018ucc ; Mondkar:2021qsf .
@@
The phenomenological implications of the semi-holographic approach, which can be formulated consistently at any level of coarse-graining, can be derived primarily from the hybrid hydrodynamic attractor surface which was studied explicitly in Mitra:2020mei ; Mitra:2022uhv in the context of the Bjorken flow by approximating the dynamical description of each subsector in terms of a Müller-Israel-Stewart (MIS) formulation Muller:1967zza ; Israel:1979wp ; Baier:2007ix . It was shown that not only on the attractor surface, but also for phenomenologically relevant flows (evolving towards the attractor surface) which can be matched to glasma-type approaches Kowalski:2007rw at relevant proper times, the bottom-up thermalization scenario Baier:2000sb is realized (i.e. it can be demonstrated analytically that the energy density of the hard sector dominates over that of the soft sector) while the full system can be described in terms of a single fluid at late time. Furthermore, the comparison of hydrodynamization of the individual subsectors leads to novel insights into small system collisions. It was shown that although the perturbative sector hydrodynamizes later in Pb-Pb type collisions compared to p-p type collisions, the soft sector can hydrodynamize earlier in p-p type collisions compared to Pb-Pb type collisions, both on the hybrid attractor surface and in phenomenologically relevant flows Mitra:2022uhv . This follows from how the hydrodynamization times scale with the total energy density in both subsectors.
@@
In this work, we investigate fundamental issues about thermalization in the semi-holographic approach. The two subsystems are isolated in their effective metrics, although they do exchange energy and momentum from the point of view of the physical background metric. In the large $N$ limit, the subsystems also have their individual entropy currents. This implies that, for generic initial conditions, the full system can relax to pseudo-equilibrium states, where each subsystem has a distinct physical temperature, without reaching global equilibrium. In fact, in Sec.˜5, we explicitly demonstrate that the latter can indeed be the case. Furthermore, in the democratic coupling scheme, it is not obvious a priori that the following are satisfied:
@@
In Kurkela:2018dku , conditions (i) and (ii) were shown to hold generally in the semi-holographic democratic coupling scheme. We review these proofs in Sec.˜3 while stating the issue of pseudo-equilibrium states explicitly and extending the concept of statistical entropy to the latter. In Sec.˜4 we prove that the condition (iii) mentioned above is indeed satisfied, i.e., the global equilibrium state, which can be consistently defined, is indeed the unique maximum entropy state of the micro-canonical ensemble.
The formulation of semi-holography allows us to study the dynamics of the full system at any level of coarse-graining. We will address the fundamental issues in thermalization by employing effective MIS-type descriptions of the two subsystems in which a microstate configuration of the full system is specified phenomenologically by the sub-system energy-momentum tensor themselves. This allows us to study thermalization by following the evolution of the coarse-grained (phenomenological) entropies of the sub-systems and the full system as elaborated in Sec. 2.2 via the example of two coupled effective kinetic theories.
In Sec. 5, we employ the coarse-grained MIS description for both sectors and demonstrate that when the average total energy density is large, a typical non-equilibrium state of the full isolated system does indeed relax to the global equilibrium state, by formulating this criterion in precise mathematical terms. Finally, in Sec. 6, we compare our results with quantum statistical approaches such as the eigenstate thermalization hypothesis, and discuss further studies which are necessary for phenomenological applications.
Figure 1: The isolated system, $\mathcal{F}$, is composed of two subsystems, $\mathcal{U}$ and $\tilde{\mathcal{U}}$, which individually satisfy their own conservation equations eq.˜4 in their respective background metric. The two subsystems are coupled and interact via deformations of their metric, given by sec.˜2.1.
@@ 2 Semi-holography
The semi-holographic framework proposes a democratic formulation Banerjee:2017ozx in which at any energy scale, the full theory is defined by both the weakly coupled perturbative sectors as well as the strongly coupled non-perturbative sector, which has a classical gravity dual in the large $N$ limit. The two sectors, otherwise isolated from each other, are allowed to interact by deforming the marginal and relevant couplings of each sector, which get promoted to local algebraic functions of the corresponding operators, and this can be implemented via an effective action Banerjee:2017ozx ; Kurkela:2018dku . This coupling scheme ensures the renormalizability of the effective action of the full system in the large $N$ limit. The democratic coupling also allows for the construction of low-energy dynamics of the full system from the effective low-energy descriptions of the individual sectors. The two sectors living in their respective effective metric backgrounds are assumed to share the same topological space so that we can describe the full system, including these sectors, using the same coordinates. Moreover, the energy-momentum tensor of the full system is conserved in the physical background metric while the subsectors are closed in their respective background metrics and are allowed to exchange energy and momentum with respect to the physical background metric. This picture is summarized in Fig.˜1.
@@ 2.1 Effective metric coupling
Consider a dynamical system $\mathcal{F}$ in a fixed metric background $g^{(B)}_{\mu\nu}$. The physical system is composed of two sub-sectors $\mathcal{U}$ and $\tilde{\mathcal{U}}$, which correspond to the perturbative and non-perturbative degrees of freedom, see Fig.˜1. The two sub-sectors live in their respective metric background $g_{\mu\nu}$ and $\tilde{g}_{\mu\nu}$. The effective action of the system that captures the deformation of the marginal/relevant coupling of the individual sectors and their interaction is Kurkela:2018dku
@@
in $d$-dimensional spacetime. S{S} and $\tilde{{S}}$ are the effective actions of the respective sectors, $\mathcal{U}$ and $\tilde{\mathcal{U}}$ and $\phi$ and $\tilde{\phi}$ denote all the matter fields in the two sub-sectors. $\gamma$ and $\gamma^{\prime}$ are the dimensionful semi-holographic coupling of mass dimension d-d whose ratio gives the dimensionless coupling $r:=-\frac{\gamma^{\prime}}{\gamma}$. The effective metrics $g_{\mu\nu}$ and $\tilde{g}_{\mu\nu}$ act as auxiliary fields and capture the marginal deformation of the respective sectors. As mentioned earlier, the deformation is taken care of by promoting the effective metrics as local algebraic functions of the operators of the sub-sectors. The algebraic equations of the effective metric are obtained by varying the action (1) with respect to the individual effective metrics. The lowest-order metric coupling equations that dictate the coupling between the sub-sectors explicitly are
@@
$T^{{\mu\nu}}$ $:=$ $-\frac{2}{\sqrt{-g^{(B)}}}\frac{\delta{S}_{tot}}{\delta g^{(B)}_{\mu\nu}}$,
@@
$\Delta K=$ - $\frac{\gamma}{2}~\Bigg(t^{\rho\alpha}\frac{\sqrt{-g}}{\sqrt{-g^{(B)}}}\Bigg)~g_{\alpha\beta}^{(B)}~\Bigg(\tilde{t}^{\beta\sigma}\frac{\sqrt{-\tilde{g}}}{\sqrt{-g^{(B)}}}\Bigg)~g^{(B)}_{\sigma\rho}$ (6)
@@
As the full action is not invariant under individual diffeomorphisms of the subsystems but only under their diagonal combination, the Ward identities of the individual subsystems imply the Ward identity for the full system.
@@
2.2 On microstates and coarse-grained entropy in semi-holography
One of the main advantanges of the democratic effective couplings described above is that microstates of the full system can be defined in the semi-holographic framework at any level of coarse-graining in the large $N$ limit. We will not pursue this explicitly here in the full quantum field theory although this follows from the action formulation (1). Assuming that both the hard (UV) and soft (IR) sectors are weakly coupled, we can use effective kinetic theories to describe each sector. Here we illustrate that we can define the microstates and their time-evolution in this case as the energy-momentum tensor and effective metric of each sub-sector is well-defined self-consistently together in terms of the microscopic variables, namely the particle distribution functions. This can be generalized to other contexts, e.g. if we replace the kinetic description of the soft (IR) sector with a holographic description in terms of a classical theory of gravity in one higher dimension. [^2] In this case, the initial condition of the holographic sector is stated in terms of gravitational initial conditions. For an illustration see Ecker:2018ucc ; Mondkar:2021qsf in which classical fields describing the UV sector is coupled with a holographic IR sector.
Consider the two sub-systems $\mathcal{U}$ and $\tilde{\mathcal{U}}$ to be gases of weakly interacting massless particles living in their respective background metrics, $g_{\mu\nu}$ and $\tilde{g}_{\mu\nu}$. Let us assume that we can employ effective kinetic theory descriptions for each of these sub-systems. For each particle in sub-system $\mathcal{U}$, the on-shell four momentum $p^{\mu}$ should satisfy $g_{\mu\nu}p^{\mu}p^{\nu}=0$ and similarly in sub-system $\tilde{\mathcal{U}}$, the on-shell four-momentum of each massless particle satisfies $\tilde{g}_{\mu\nu}\tilde{p}^{\mu}\tilde{p}^{\nu}=0$. Clearly, we can solve the on-shell conditions to obtain the on-shell effective energies $p^{0}(x)$ and $\tilde{p}^{0}(x)$ as functions of $p^{i}$ and $g_{\mu\nu}(x)$, and $\tilde{p}^{i}$ and $\tilde{g}_{\mu\nu}(x)$, respectively. The microstates of $\mathcal{U}$ and $\tilde{\mathcal{U}}$ are specified by the particle distribution functions $f(x,p^{i})$ and $\tilde{f}(x,\tilde{p}^{i})$, respectively at a given point of time $x^{0}$.
The distribution functions allow us to compute the macroscopic quantities, particularly the energy-momentum tensor $t^{\mu\nu}$ and $\tilde{t}^{\mu\nu}$ for each sector,
$t^{\mu\nu}(x)=\int\frac{d^{3}p}{(2\pi)^{3}\sqrt{-g(x)}}\frac{p^{\mu}(x)p^{\nu}(x)}{p^{0}(x)}f(x,p^{i}),$
$\tilde{t}^{\mu\nu}(x)=\int\frac{d^{3}\tilde{p}}{(2\pi)^{3}\sqrt{-\tilde{g}(x)}}\frac{\tilde{p}^{\mu}(x)\tilde{p}^{\nu}(x)}{\tilde{p}^{0}(x)}\tilde{f}(x,\tilde{p}^{i}).$ (8)
Note that $t^{\mu\nu}(x)$ and $\tilde{t}^{\mu\nu}(x)$ should be obtained self-consistently along with the effective metrics $g_{\mu\nu}(x)$ and $\tilde{g}_{\mu\nu}(x)$ by solving the algebraic coupling equations (2.1). All of these can be determined locally given the microscopic variables, $f(x,p^{i})$ and $\tilde{f}(x,\tilde{p}^{i})$.
It could be useful to see how one can define the kinetic equations for each sub-sector in the RTA (relaxation time approximation). First, we need to define local equilibrium particle distributions via matching conditions (see e.g. Romatschke:2015gic ; Bajec:2024jez ; Bajec:2025dqm ). To do so, we construct the flow vectors $u^{\mu}$ and $\tilde{u}^{\mu}$ in the Landau frame so that these are time-like eigenvectors of $t^{\mu}_{\,\nu}$ and $\tilde{t}^{\mu}_{\,\nu}$, respectively, and they satisfy $u^{\mu}g_{\mu\nu}u^{\nu}=\tilde{u}^{\mu}\tilde{g}_{\mu\nu}\tilde{u}^{\nu}=-1$. The energy densities of $\mathcal{U}$ and $\tilde{\mathcal{U}}$ are
$\varepsilon=\int\frac{d^{3}p}{(2\pi)^{3}\sqrt{-g}}\frac{(u_{\mu}(x)p^{\mu}(x))^{2}}{p^{0}(x)}f(x,p^{i}),$
$\tilde{\varepsilon}=\int\frac{d^{3}\tilde{p}}{(2\pi)^{3}\sqrt{-g}(x)}\frac{(u_{\mu}(x)p^{\mu}(x))^{2}}{p^{0}(x)}f(x,p^{i}),$ (9)
respectively. Finally, we can define the local equilibrium particle distributions
$f_{eq}=\frac{1}{e^{-p^{\mu}u^{\nu}g_{\mu\nu}/T_{1}}-1},\,\,\,\,\,\,\,\tilde{f}_{eq}=\frac{1}{e^{-\tilde{p}^{\mu}\tilde{u}^{\nu}\tilde{g}_{\mu\nu}/T_{2}}-1}.$ (10)
where $T_{1}$ and $T_{2}$ are defined from $\varepsilon(T_{1})$ and $\tilde{\varepsilon}(T_{2})$, respectively, via the equations of states of the respective subsystems. The evolution of the one-particles distribution function is given by the RTA Boltzmann equation
$p^{\mu}\partial_{\mu}f+p^{\mu}p^{\nu}\Gamma^{i}_{\mu\nu}\frac{\partial f}{\partial p_{i}}=\frac{p^{\mu}u_{\mu}}{\tau_{R}}(f-f_{eq})$ (11)
for $\mathcal{U}$ with $\tau_{R}=C_{\tau}T_{1}^{-1}$ and $C_{\tau}$ a constant. Similarly, we can construct the RTA Boltzmann equation for $\tilde{\mathcal{U}}$ with $\tilde{\tau}_{R}=\tilde{C}_{\tau}T_{2}^{-1}$ and $\tilde{C}_{\tau}$ a constant.
It is not hard to see that the full dynamics can be solved self-consistently with initial conditions where $f(x,p^{i})$ and $\tilde{f}(x,\tilde{p}^{i})$ are given at any point of time. We will explicitly describe an algorithm in a simplified context in this paper where each subsector is further simplified with a BRSSS description as in Mitra:2022uhv . The interested reader can study another illustration of semi-holographic dynamics in which a quantum harmonic oscillator is coupled to a quantum dot which admits a holographic dual description in terms of a two-dimensional gravitational theory Kibe:2023ixa .
Our illustration with coupled effective kinetic theories also demonstrates that just like the full dynamics we can also define coarse-grained entropies of the subsystems at any level of coarse-graining. In case of the full quantum theory, the coarse-grained entropy of a state is defined as the von-Neuman entropy of the density matrix which has the maximum von-Neuman entropy in the set of all density matrices which reproduce the mean values of simple observables in the given state. In the context of the effective kinetic descriptions, we can define the entropy currents of the (bosonic) sub-systems via
$s^{\mu}=g_{d}\int\frac{d^{3}p}{(2\pi)^{3}\sqrt{-g}}\frac{p^{\mu}}{p^{0}}[f\ln f-(1+f)\ln(1+f)],$
[…]
where $g_{d}$ and $\tilde{g}_{d}$ are the respective degeneracy factors. If the kinetic description of the soft sub-system is replaced by a holographic description, a phenomenological entropy current can be derived from the apparent horizon (see e.g. Booth:2010kr ; Jansen:2020ign ).
To understand fundamental issues about thermalization, we will employ effective BRSSS descriptions of the perturbative and non-perturbative subsystems assuming them to be conformal (note that the effective metric coupling breaks conformality). In this context the microstates of the sub-systems are specified by their energy-momentum tensors ($t^{\mu}_{\,\nu}$ and $\tilde{t}^{\mu}_{\,\nu}$) themselves instead of fine-grained variables such as the particle distribution functions. In this case, the microstates and the entropies are defined phenomenologically. This simplification will allow us to study the evolution of the coarse-grained (phenomenological) entropies of the sub-systems and the full system, and thus thermalization.
@@ 3.1 The global equilibrium
[…]
@@
[…]
@@ 3.2 Proof of thermodynamic/statistical consistencies
Both thermodynamic and statistical consistencies follow from the general form of the full energy-momentum tensor. The latter given by (2.1) and (6) with one contravariant and one covariant index indicates that the terms in $\Delta K$ are always diagonal and appear with opposite signs for $T^{0}_{\,0}$ and $T^{i}_{\,i}$, i.e., for $\mathcal{E}$ and $\mathcal{P}$. This leads to the identity
@@ 3.3 The persistence of pseudo-equilibrium
However, there is another problem of dynamical origin. More generally, if the two subsystems have entropy currents $s_{1}^{\mu}$ and $s_{2}^{\mu}$ (which in equilibrium take the forms $s_{1}^{\mu}=s_{1}u^{\mu}$ and $s_{2}^{\mu}=s_{2}\tilde{u}^{\mu}$) that satisfy $\nabla_{\mu}s_{1}^{\mu}\geq 0$ and $\tilde{\nabla}_{\mu}s_{2}^{\mu}\geq 0$ in their respective backgrounds, then it is easy to see that the total entropy current $\mathcal{S}^{\mu}$ can be defined as